# CropVLM: Learning to Zoom for Fine-Grained Vision-Language Perception

### v1

@@ 1 Introduction
Recent VLMs have demonstrated impressive capabilities in understanding and reasoning over visual contents [alayrac2022flamingo, liu2023visual, chen2024internvl]. Still, despite their success, these models face significant limitations when confronted with tasks that require fine-grained visual perception, such as document analysis, scene-text recognition, or detailed object identification. A primary constraint lies in the input resolution, as most mainstream VLMs leverage pre-trained vision encoders that can take as input images at relatively low resolutions – e.g., 224×224 [radford2021learning] or 336×336 pixels [liu2023visual] – causing crucial fine details to become indiscernible.
@@
For instance, the LLaVA-1.5 [liu2023visual] model, with an input resolution of 336x336 pixels, is effective for general visual reasoning, but it struggles significantly when processing small text or when performing detailed visual analysis [liu2024ocrbench]. A straightforward solution involves increasing the input resolution uniformly. However, this introduces a prohibitive computational burden with VLMs based on the Transformer architecture. Recent work by shi2025scaling or cai2024matryoshka demonstrates that even advanced models use only a small number of image tokens to answer most requests, suggesting that uniform high-resolution processing is inefficient and unnecessary. Alternative approaches have attempted to address these limitations through architectural modifications [cai2024matryoshka, shi2024we] or specialized fine-tuning [shao2024visual]. While effective in controlled settings, these methods often require extensive model retraining, risking catastrophic forgetting and poor generalization to out-of-domain scenarios. Furthermore, these strategies are typically not applicable to proprietary models, given their closed weights.
@@
SEAL [wu2024v] LLM-Guided Search LLM-Only Multiple
SEMCLIP [li2025semantic] Semantic Embedding-Guided Single
ViCrop [zhang2025mllms] Semantic Attention-Heuristic Single
Visual-CoT [shao2024visual] Manual Human & Synthetic SFT Single
VisRL [chen2025visrl] Preference Model Validated DPO Single
UV-CoT [zhao2025unsupervised] Preference Model Validated DPO Single
Visual-RFT [liu2025visual] Preference Human GRPO Single
DeepEyes [zheng2025deepeyes] Preference GRPO Multiple
Chain-of-Focus [zhang2025chain] Preference GRPO Multiple
Mini-o3 [lai2025mini] Preference GRPO Multiple
@@
The fundamental challenge lies in the inability of current VLMs to dynamically adjust the visual focus across different spatial regions based on the task at hand, even when guided by detailed textual prompts [zhao2025unsupervised]. This limitation highlights the critical need for a flexible approach that can adaptively allocate computational resources to the most relevant parts of an input image.
@@
Targeting the aforementioned advantages, this paper introduces CropVLM [^1] The code is made available in https://github.com/miguelscarv/cropvlm as a reinforcement learning-based approach that enables VLMs to dynamically ”zoom in” on relevant image regions, without requiring ground-truth bounding boxes as the supervision signal, and without a separate evaluator model guiding the training. Our method enhances existing VLMs with a lightweight cropping network (i.e., with only 256M parameters), which identifies task-relevant regions for finer-detail image processing. We leverage reinforcement learning to train the region selection model, eliminating the need for expensive human annotations in the form of bounding boxes. Our approach functions as a modular component that can be paired with both open-source and proprietary VLMs, without requiring modifications over the target model.
@@ 2 Related Work
Despite their impressive capabilities, VLMs exhibit notable limitations in visual perception, particularly in tasks requiring fine-grained understanding [tong2024eyes]. These challenges are especially pronounced in tasks such as document understanding, where precise visual parsing is critical. While VLMs often excel at textual reasoning, their ability to interpret fine grained details, e.g. such as small text regions [liu2024ocrbench] or small objects, remains suboptimal, motivating recent efforts to improve visual grounding.
@@
A common approach to improving visual perception is Supervised Fine-Tuning (SFT) on task-specific datasets. Alternatively, other approaches focus on architectural improvements. For example, Matryoshka (M3) [cai2024matryoshka] uses progressive training to enhance high-resolution reasoning, and S2 [shi2024we] adapts models across visual granularities. The PS3 approach [shi2025scaling] scales pre-training to 4K resolution by contrasting local image regions with detailed captions. Region-aware strategies such as Visual-CoT [shao2024visual] use bounding-box supervision to help models focus on task-relevant regions, but often rely on expensive annotations or narrow training distributions.
@@
To make visual grounding more adaptive, intention-driven methods dynamically prioritize image regions based on user queries. These methods fall into three broad categories: (1) LLM-guided querying, such as SEAL [wu2024v], SEMCLIP [li2025semantic], or SoM-LLaVA [yan2024list], which guide region selection using search, textual semantics, or visual tags; (2) cropping-based preprocessing, as seen in Visual-CoT [shao2024visual] and ViCrop [zhang2025mllms], which crop inputs before answering the request; and (3) preference-based refinement, such as VisRL [chen2025visrl], which optimize cropping through reward models or preference feedback. While these techniques improve relevance and efficiency, they often require expensive human or synthetic annotations. In contrast, the proposed CropVLM method simplifies the process of learning how to focus without requiring explicit region supervision.
@@
Preference optimization has emerged as a scalable framework for aligning models with human intent [ouyang2022training]. Visual-CoT [shao2024visual] applies supervised fine-tuning using annotated regions, while UV-CoT [zhao2025unsupervised] and VisRL [chen2025visrl] remove this bottleneck by generating synthetic region preferences for Direct Preference Optimization (DPO) via textual feedback from general VLMs. The Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] approach offers advantages over DPO in simplified reward modeling and stability. For example, Visual-RFT [liu2025visual] applies GRPO to perception and image classification tasks, using verifiable rewards based on ground-truth bounding boxes and labels. Recent works such as DeepEyes [zheng2025deepeyes], Chain-of-Focus [zhang2025chain], and Mini-o3 [lai2025mini] extend this line of research through multi-turn region-of-interest reasoning. While powerful, these approaches require multiple reasoning steps at inference, leading to substantially higher test-time cost. Moreover, it remains unclear how to identify the most informative crop within a model’s reasoning trace to effectively assist a target VLM in answering a query. Our work instead applies GRPO to single-pass visual cropping, achieving alignment without manual or synthetic bounding boxes and enabling scalable optimization, even for closed-weight or out-of-domain settings.
@@ 3 Methods
This section introduces the CropVLM training method. In brief, CropVLM is a Vision-Language Model (VLM) trained to produce a single bounding box per request-image pair, aiming to improve downstream performance with a target (separate) VLM. The bounding box is generated as text in the format $x_{1},y_{1},x_{2},y_{2}]$, where $(x_{1},y_{1})$ denotes the top-left corner and $(x_{2},y_{2})$ the bottom-right corner of the crop, and where each coordinate corresponds to a percentage of the width and height of the image, so as to generalize to differently sized images.
@@
The specific reinforcement learning algorithm employed in our training procedure is Group Relative Policy Optimization (GRPO) [shao2024deepseekmath], i.e. a variant of Proximal Policy Optimization (PPO) [schulman2017proximal] which simplifies training by avoiding the need for a value model. This algorithm allows us to optimize the bounding box generation policy in an unsupervised manner. Unlike other reinforcement learning methods such as PPO, that require a separate critic model to evaluate policy performance, GRPO directly compares groups of candidate responses. For a given input query $q$, the algorithm generates $G$ distinct responses $\{o_{1},o_{2},\ldots,o_{G}\}$ from the current policy $\pi_{\theta_{\text{old}}}$. It then evaluates these responses and assigns rewards $\{r_{1},r_{2},\ldots,r_{G}\}$ based on their quality. GRPO normalizes the rewards by computing their mean and standard deviation, this way determining the relative quality $A_{i}$ of each response:
@@
Previous work has shown that concatenating the full image with a relevant image crop can improve VQA accuracy [zhang2025mllms]. We apply a similar strategy by combining the full image and the crop, afterwards passing these inputs to a target VLM from which a reward signal is computed, as shown in Figure 2.
@@
  • Likelihood-Based Reward: The most common ground-truth answer for the VQA sample is extracted from the dataset. The VLM then computes the log-likelihood of this answer given the input. This formulation allows for faster, single-pass reward estimation, without requiring autoregressive generation. Formally, we define the likelihood-based reward as follows:
@@
    where $p_{\theta}(a^{*}_{t}|I_{o},I_{c},q,a^{*}_{<t})$ is the probability assigned by the reward VLM to the $t$-th token of the ground-truth answer $a^{*}$, given the preceding tokens $a^{*}_{<t}$, the original image $I_{o}$, the cropped image $I_{c}$, and the question $q$.
@@
Our reward formulations offer key advantages over existing approaches. Unlike VisRL [chen2025visrl], which uses general-purpose VLMs for textual rewards without specialized training, our method directly measures performance using quantifiable metrics aligned with the downstream task. Additionally, our GRPO-based approach achieves higher data efficiency by utilizing all $G$ sampled bounding boxes during training, rather than retaining only two responses per query, as in DPO-based methods [zhao2025unsupervised]. By grounding rewards in objective measures rather than model-generated judgments, our approach also reduces reward hacking and ensures that the improvements indeed correlate with genuine performance gains.
@@ 4 Experiments
4.1 Experimental Setup
For all subsequent experiments, we select the 256M SmolVLM Instruct model [marafioti2025smolvlm] to create CropVLM cropping networks, as this model can process images with varying resolutions and is small in size.
Datasets.
Fine-grained high-resolution image understanding is essential for tasks such as scene-text recognition and document understanding. These tasks often involve images with only a single and relatively small zone of interest per query. For these reasons, we selected the following training datasets: TextVQA [singh2019towards] and ST-VQA [biten2019scene] for scene-text understanding, DocVQA [mathew2021docvqa] for document analysis, and InfographicsVQA [mathew2022infographicvqa] for understanding graphs and infographics. This mixture results in a total of 124k image-question pairs. Since SmolVLM is not originally capable of generating valid bounding boxes in the format we described in Section 3, we randomly split the above training mixture into 2 equally-sized datasets, where one is used to create a synthetic seed bounding box dataset to finetune SmolVLM on, using an initial step of Supervised Fine-Tuning (SFT), and the other is used to do GRPO, after the aforementioned SFT step. To create the seed bounding box dataset we prompted the Qwen 2.5-VL 7B Instruct [bai2025qwen2] model with the instruction ”Outline the region in the image that would help answer the following question: {QUESTION} Output the coordinates in JSON format.”, to obtain a synthetic bounding box dataset. Our goal is to give the model the basic ability of generating bounding boxes in the intended format. Since the Qwen 2.5-VL model generates absolute coordinate bounding boxes, an extra normalization step is performed to comply with the previously designed format. Additionally, upon inspection of the generated bounding boxes, we observed that the ones with smaller relative areas tend to not include the correct area of interest within the image. To address this issue, we expand the generated bounding boxes according to their relative area percentiles, using the expansion factors reported in Table 3.
Relative Area (%) Percentile Expansion Factor
<< 0.16% Below 20th 45x
0.16% - 0.38% 20th - 40th 10x
0.38% - 0.91% 40th - 60th 4x
0.91% - 3.51% 60th - 80th 2x
>> 3.51% Above 80th 1x (No expansion)
The use of external bounding boxes for the SFT stage is ablated in Appendix E, given in the supplementary material, showing that CropVLM achieves similar performance even when initialized through a simpler procedure that does not involve an extra model.
Training Stages.
The Supervised Finetuning (SFT) stage trained 3 separate SmolVLM models at 3 different input resolutions, i.e. 512x512, 1024x1024 and 2048x2048 pixels, on the synthetic bounding box dataset. These SmolVLM models can produce rough bounding box extimates in response to the prompt: ”{QUESTION} Outline the region in the image that would help answer this question.”. Then, the GRPO stage improved each of these models at the aforementioned resolutions on the remaining half of the dataset. In the GRPO stage, the generated bounding boxes are evaluated using a reward model that corresponds to a base 256M SmolVLM Instruct model with an input resolution of 512x512 pixels, motivated by efficiency. Using a lightweight model as the target VLM, from which rewards are derived, ensures fast evaluation while still being sufficiently accurate to provide meaningful learning signals, particularly because of group-normalized rewards. The experiments in Section 4.2.2 show that CropVLM, framed with this reward model, can also be used to improve the results of larger VLMs. As described in Section 3, we experimented with 2 reward types, i.e. accuracy and likelihood-based, in addition to a bounding box format validation reward. The bounding box validation reward takes the value of 1 if the bounding box is valid and 0 otherwise, when training with likelihood-based rewards, and 0.25 if valid and 0 otherwise, for the accuracy-based rewards. This difference in scaling was necessary since the two types of rewards operate on different scales. The accuracy reward used each dataset’s official metrics: VQA accuracy for TextVQA, and ANLS (Average Normalized Levenshtein Similarity) for DocVQA, ST-VQA, and InfographicVQA. The likelihood reward was calculated using the log-likelihood over the ground truth response tokens. In both cases, the prompt used in the reward model was the following: ”{QUESTION} Give a very brief answer.”.
Hyperparameters.
For all training runs, we employed LoRA [hu2022lora] over the SmolVLM model with consistent parameters: rank 128, alpha 256, and dropout of 0.05. We maintained a batch size of 16 across all tests and used a cosine decay learning rate scheduler. All models were trained for a single epoch. Stage-specific hyperparameters varied as follows: For the SFT stage, we used a learning rate of 5e-5 and maximum gradient norm of 1.0. For the GRPO stage, we reduced these values to 5e-6 and 0.1, respectively. Additional GRPO-specific hyperparameters included a group size of 6, generation temperature of 0.8, and beta of 0.01. These conservative hyperparameters, namely the small group size in GRPO, together with the use of limited reinforcement learning data, and also small model sizes, were necessary due to resource constraints. Specifically, all our models were trained on a single A100 GPU. Using these hyperparameters, the SFT stage required approximately 3 GPU hours while the GRPO stage required 24 GPU hours for our largest resolution CropVLM model at 2048px. As a result, our reported results likely represent a lower bound on the full potential of the CropVLM method under training regimes that involve more resources.
4.2 Experimental Results
To assess the performance of our method, we evaluated on splits of the same datasets used during training. Specifically, for document and infographics understanding, we use the validation splits of DocVQA and InfographicsVQA. For scene-text analysis, we use the validation split of TextVQA and the test split of ST-VQA. To assess zero-shot high-resolution general vision understanding, we evaluate on the V* [wu2024v] benchmark and also on the recent HR-Bench 4k and 8k [wang2025divide] benchmarks. For each dataset, we employ the official evaluation metric: VQA accuracy for TextVQA, V* and HR-Bench 4k and 8k, and ANLS for DocVQA, ST-VQA and InfographicVQA. Appendices B, D, and H respectively contain statistics for bounding boxes generated by our method, ablations on the use of the full image, and qualitative examples obtained with CropVLM.
4.2.1 Evaluating CropVLM Paired with SmolVLM

### v2

@@ 1 Introduction
Recent VLMs have demonstrated impressive capabilities in understanding and reasoning over visual contents [1, 11, 6]. Still, despite their success, these models face significant limitations when confronted with tasks that require fine-grained visual perception, such as document analysis, scene-text recognition, or detailed object identification. A primary constraint lies in the input resolution, as most mainstream VLMs leverage pre-trained vision encoders that can take as input images at relatively low resolutions – e.g., 224×224 [18] or 336×336 pixels [11] – causing crucial fine details to become indiscernible.
@@
For instance, the LLaVA-1.5 [11] model, with an input resolution of 336x336 pixels, is effective for general visual reasoning, but it struggles significantly when processing small text or when performing detailed visual analysis [12]. A straightforward solution involves increasing the input resolution uniformly. However, this introduces a prohibitive computational burden with VLMs based on the Transformer architecture. Recent work by Shi et al. [22] or Cai et al. [4] demonstrates that even advanced models use only a small number of image tokens to answer most requests, suggesting that uniform high-resolution processing is inefficient and unnecessary. Alternative approaches have attempted to address these limitations through architectural modifications [4, 23] or specialized fine-tuning [20]. While effective in controlled settings, these methods often require extensive model retraining, risking catastrophic forgetting and poor generalization to out-of-domain scenarios. Furthermore, these strategies are typically not applicable to proprietary models, given their closed weights.
@@
SEAL [27] LLM-Guided Search LLM-Only Multiple
SEMCLIP [10] Semantic Embedding-Guided Single
ViCrop [29] Semantic Attention-Heuristic Single
Visual-CoT [20] Manual Human & Synthetic SFT Single
VisRL [5] Preference Model Validated DPO Single
UV-CoT [31] Preference Model Validated DPO Single
Visual-RFT [13] Preference Human GRPO Single
DeepEyes [32] Preference GRPO Multiple
Chain-of-Focus [30] Preference GRPO Multiple
Mini-o3 [9] Preference GRPO Multiple
@@
The fundamental challenge lies in the inability of current VLMs to dynamically adjust the visual focus across different spatial regions based on the task at hand, even when guided by detailed textual prompts [31]. This limitation highlights the critical need for a flexible approach that can adaptively allocate computational resources to the most relevant parts of an input image.
@@
Targeting the aforementioned advantages, this paper introduces CropVLM [^1] The code is made available at https://github.com/miguelscarv/cropvlm as a reinforcement learning-based approach that enables VLMs to dynamically ”zoom in” on relevant image regions, without requiring ground-truth bounding boxes as the supervision signal, and without a separate evaluator model guiding the training. Our method enhances existing VLMs with a lightweight cropping network (i.e., with only 256M parameters), which identifies task-relevant regions for finer-detail image processing. We leverage reinforcement learning to train the region selection model, eliminating the need for expensive human annotations in the form of bounding boxes. Our approach functions as a modular component that can be paired with both open-source and proprietary VLMs, without requiring modifications over the target model.
@@ 2 Related Work
Despite their impressive capabilities, VLMs exhibit notable limitations in visual perception, particularly in tasks requiring fine-grained understanding [25]. These challenges are especially pronounced in tasks such as document understanding, where precise visual parsing is critical. While VLMs often excel at textual reasoning, their ability to interpret fine grained details, e.g. such as small text regions [12] or small objects, remains suboptimal, motivating recent efforts to improve visual grounding.
@@
A common approach to improving visual perception is Supervised Fine-Tuning (SFT) on task-specific datasets. Alternatively, other approaches focus on architectural improvements. For example, Matryoshka (M3) [4] uses progressive training to enhance high-resolution reasoning, and S2 [23] adapts models across visual granularities. The PS3 approach [22] scales pre-training to 4K resolution by contrasting local image regions with detailed captions. Region-aware strategies such as Visual-CoT [20] use bounding-box supervision to help models focus on task-relevant regions, but often rely on expensive annotations or narrow training distributions.
@@
To make visual grounding more adaptive, intention-driven methods dynamically prioritize image regions based on user queries. These methods fall into three broad categories: (1) LLM-guided querying, such as SEAL [27], SEMCLIP [10], or SoM-LLaVA [28], which guide region selection using search, textual semantics, or visual tags; (2) cropping-based preprocessing, as seen in Visual-CoT [20] and ViCrop [29], which crop inputs before answering the request; and (3) preference-based refinement, such as VisRL [5], which optimize cropping through reward models or preference feedback. While these techniques improve relevance and efficiency, they often require expensive human or synthetic annotations. In contrast, the proposed CropVLM method simplifies the process of learning how to focus without requiring explicit region supervision.
@@
Preference optimization has emerged as a scalable framework for aligning models with human intent [17]. Visual-CoT [20] applies supervised fine-tuning using annotated regions, while UV-CoT [31] and VisRL [5] remove this bottleneck by generating synthetic region preferences for Direct Preference Optimization (DPO) via textual feedback from general VLMs. The Group Relative Policy Optimization (GRPO) [21] approach offers advantages over DPO in simplified reward modeling and stability. For example, Visual-RFT [13] applies GRPO to visual perception and image classification tasks, using verifiable rewards based on ground-truth bounding boxes and labels. Recent studies such as DeepEyes [32], Chain-of-Focus [30], or Mini-o3 [9] extend this line of research through multi-turn region-of-interest reasoning. While powerful, these approaches require multiple reasoning steps at inference, leading to substantially higher test-time cost. Moreover, it remains unclear how to identify the most informative crop within a model’s reasoning trace, to effectively assist a target VLM in answering a query. Our work instead applies GRPO to single-pass visual cropping, achieving alignment without manual or synthetic bounding boxes, and enabling scalable optimization, even for closed-weight or out-of-domain settings.
@@ 3 Methods
This section introduces the CropVLM training method. In brief, CropVLM is a relatively small Vision-Language Model (VLM) trained to produce a single bounding box per request-image pair, aiming to improve downstream performance with a target (separate) VLM. The bounding box is generated as text in the format $x_{1},y_{1},x_{2},y_{2}]$, where $(x_{1},y_{1})$ denotes the top-left corner and $(x_{2},y_{2})$ the bottom-right corner of the crop, and where each coordinate corresponds to a percentage of the width and height of the image, so as to generalize to differently sized images.
@@
The specific reinforcement learning algorithm employed in our training procedure is Group Relative Policy Optimization (GRPO) [21], i.e. a variant of Proximal Policy Optimization (PPO) [19] which simplifies training by avoiding the need for a value model. This algorithm allows us to optimize the bounding box generation policy in an unsupervised manner. Unlike other reinforcement learning methods such as PPO, that require a separate critic model to evaluate policy performance, GRPO directly compares groups of candidate responses. For a given input query $q$, the algorithm generates $G$ distinct responses $\{o_{1},o_{2},\ldots,o_{G}\}$ from the current policy $\pi_{\theta_{\text{old}}}$. It then evaluates these responses and assigns rewards $\{r_{1},r_{2},\ldots,r_{G}\}$ based on their quality. GRPO normalizes the rewards by computing their mean and standard deviation, this way determining the relative quality $A_{i}$ of each response:
@@
Previous work has shown that concatenating the full image with a relevant image crop can improve VQA accuracy [29]. We apply a similar strategy by combining the full image and the crop, afterwards passing these inputs to a target VLM from which a reward signal is computed, as shown in Figure 2.
@@
  • Likelihood-Based Reward: The most common ground-truth answer for the request is extracted from the dataset. The VLM then computes the log-likelihood of this answer given the corresponding input. This formulation allows for faster, single-pass reward estimation, without requiring autoregressive generation. Formally, we define the likelihood-based reward as follows:
@@
    where $p_{\theta}(a^{*}_{t}|I_{o},I_{c},q,a^{*}_{<t})$ is the probability assigned by the reward VLM to the $t$-th token of the ground-truth answer $a^{*}$, given the preceding tokens $a^{*}_{<t}$, the original image $I_{o}$, the cropped image $I_{c}$, and the request $q$.
@@
Our reward formulations offer key advantages over existing approaches. Unlike VisRL [5], which uses general-purpose VLMs for textual rewards without specialized training, our method directly measures performance using quantifiable metrics aligned with the downstream task. Additionally, our GRPO-based approach achieves higher data efficiency by utilizing all $G$ sampled bounding boxes during training, rather than retaining only two responses per query, as in DPO-based methods [31]. By grounding rewards in objective measures rather than model-generated judgments, our approach also reduces reward hacking and ensures that the improvements indeed correlate with genuine performance gains.
@@ 4 Experiments
4.1 Experimental Setup
For all subsequent experiments, we select the 256M SmolVLM Instruct model [14] to create CropVLM cropping networks, as this model can process images with varying resolutions and is small in size.
Datasets.
Fine-grained high-resolution image understanding is essential for tasks such as scene-text recognition and document understanding. These tasks often involve images with only a single and relatively small zone of interest per query. For these reasons, we selected the following training datasets: TextVQA [24] and ST-VQA [3] for scene-text understanding, DocVQA [16] for document analysis, and InfographicsVQA [15] for understanding graphs and infographics. This mixture results in a total of 124k image-question pairs. Since SmolVLM is not originally capable of generating valid bounding boxes in the format we described in Section 3, we randomly split the above training mixture into 2 equally-sized datasets, where one is used to create a synthetic seed bounding box dataset to finetune SmolVLM on, using an initial step of Supervised Fine-Tuning (SFT), and the other is used to do GRPO, after the aforementioned SFT step. To create the seed bounding box dataset we prompted the Qwen 2.5-VL 7B Instruct [2] model with the instruction ”Outline the region in the image that would help answer the following question: {QUESTION} Output the coordinates in JSON format.”, to obtain a synthetic bounding box dataset. Our goal is to give the model the basic ability of generating bounding boxes in the intended format. Since the Qwen 2.5-VL model generates absolute coordinate bounding boxes, an extra normalization step is performed to comply with the previously designed format. Additionally, upon inspection of the generated bounding boxes, we observed that the ones with smaller relative areas tend to not include the correct area of interest within the image. To address this issue, we expand the generated bounding boxes according to their relative area percentiles, using the expansion factors reported in Table 2.
Relative Area (%) Percentile Expansion Factor
<< 0.16% Below 20th 45x
0.16% - 0.38% 20th - 40th 10x
0.38% - 0.91% 40th - 60th 4x
0.91% - 3.51% 60th - 80th 2x
>> 3.51% Above 80th 1x (No expansion)
The use of external bounding boxes for the SFT stage is ablated in Appendix E, showing that CropVLM achieves similar performance even when initialized through a simpler procedure that does not involve an extra model.
Training Stages.
The Supervised Finetuning (SFT) stage trained 3 separate SmolVLM models at 3 different input resolutions, i.e. 512x512, 1024x1024 and 2048x2048 pixels, on the synthetic bounding box dataset. These SmolVLM models can produce rough bounding box extimates in response to the prompt: ”{QUESTION} Outline the region in the image that would help answer this question.”. Then, the GRPO stage improved each of these models at the aforementioned resolutions on the remaining half of the dataset. In the GRPO stage, the generated bounding boxes are evaluated using a reward model that corresponds to a base 256M SmolVLM Instruct model with an input resolution of 512x512 pixels, motivated by computational efficiency. Using a lightweight model as the target VLM, from which rewards are derived, ensures fast evaluation while still being sufficiently accurate to provide meaningful learning signals, particularly because of group-normalized rewards. The experiments in Section 4 show that CropVLM, framed with this reward model, can also be used to improve the results of larger VLMs. As described in Section 3, we experimented with 2 reward types, i.e. accuracy and likelihood-based, in addition to a bounding box format validation reward. The bounding box validation reward takes the value of 1 if the bounding box is valid and 0 otherwise, when training with likelihood-based rewards, and 0.25 if valid and 0 otherwise, for the accuracy-based rewards. This difference in scaling was necessary since the two types of rewards operate on different scales. The accuracy reward used each dataset’s official metrics: VQA accuracy for TextVQA, and ANLS (Average Normalized Levenshtein Similarity) for DocVQA, ST-VQA, and InfographicVQA. The likelihood reward was calculated using the log-likelihood over the ground truth response tokens. In both cases, the prompt used in the reward model was the following: ”{QUESTION} Give a very brief answer.”.
Hyperparameters.
For all training runs, we employed LoRA [8] over the SmolVLM model with consistent parameters: rank 128, alpha 256, and dropout of 0.05. We maintained a batch size of 16 across all tests and used a cosine decay learning rate scheduler. All models were trained for a single epoch. Stage-specific hyperparameters varied as follows: For the SFT stage, we used a learning rate of 5e-5 and maximum gradient norm of 1.0. For the GRPO stage, we reduced these values to 5e-6 and 0.1, respectively. Additional GRPO-specific hyperparameters included a group size of 6, generation temperature of 0.8, and beta of 0.01. These conservative hyperparameters, namely the small group size in GRPO, together with the use of limited reinforcement learning data, and also small model sizes, were necessary due to resource constraints. Specifically, all our models were trained on a single A100 GPU. Using these hyperparameters, the SFT stage required approximately 3 GPU hours while the GRPO stage required 24 GPU hours for our largest resolution CropVLM model at 2048px. As a result, our reported results likely represent a lower bound on the full potential of the CropVLM method under training regimes that involve more resources.
4.2 Experimental Results
To assess the performance of our method, we evaluated on splits of the same datasets used during training. Specifically, for document and infographics understanding, we use the validation splits of DocVQA and InfographicsVQA. For scene-text analysis, we use the validation split of TextVQA and the test split of ST-VQA. To assess zero-shot high-resolution general vision understanding, we evaluate on the V* [27] benchmark and also on the recent HR-Bench 4k and 8k [26] benchmarks. For each dataset, we employ the official evaluation metric: VQA accuracy for TextVQA, V* and HR-Bench 4k and 8k, and ANLS for DocVQA, ST-VQA and InfographicVQA. Appendices B, D, and H respectively contain statistics for bounding boxes generated by our method, ablations on the use of the full image, and qualitative examples obtained with CropVLM.
4.2.1 Evaluating CropVLM Paired with SmolVLM
@@
Post-SFT Performance.
Post-GRPO Performance.
Applying GRPO results in substantial improvements across all resolutions and reward types. The 512×512 model has the largest relative gains, where cropping helps compensate for the model’s lower perceptual capacity and visual fragmentation. Log-likelihood rewards generally outperform accuracy-based ones, likely due to the fact that they can provide more nuanced feedback during training. This option virtually eliminates examples where the reward is the same within a group, which in turn leads to more examples effectively contributing to model weight updates.
Key Insights.
4.2.2 Evaluating CropVLM with Other VLMs